# Proactive Service Assurance in 5G and B5G Networks: A Closed-Loop Algorithm for End-to-End Network Slicing

### v3

@@ V Evaluation
This section will provide an overview of our experimental setup and showcase our service assurance algorithm designed to support 5G networks with network slicing, and encompassing diverse slice categories.
V-A 5G network slicing environment
A simulation at the packet level has been developed utilizing [37] to replicate a 5G network environment that includes support for slicing features. The 5G E2E network slicing simulation includes the construction of an initial configuration comprising four distinct slices, including uRLLC (video gaming), mMTC (IoT), eMBB (HD video), and an in between application service like VoIP. Each slice has different service assurance requirements and resource demands. Fig. 5 demonstrates the implementation of our 5G network, which is reinforced by an isolated E2E network slicing mechanism that leverages virtualization technology. The 5G CN enables the construction of VNFs in a dynamic manner, as displayed by the User Plane Function (UPF) in our experimentation. This enables a single slice to accommodate either a singular VNF or a group of VNFs supported by a load balancer. Hence, the CN has the capability of facilitating the scaling of VNFs in both the vertical and horizontal dimensions. To balance traffic between VNF instances within a slice, a weighted round-robin load balancing algorithm [38] was integrated into the network slice manager. In order to deal with the guaranteed bit rate requirements in network slicing, we have implemented the Hierarchical Token Bucket (HTB) queue [39] in the router, which helps to isolate virtual network links in both TN and CN, and optimize resources.
@@
Slice Transport type Total Scale factor Network direction
@@
TABLE III: Mapping from open data to the slicing devices.
@@
The simulation was configured to generate traffic patterns that are representative of real-world networks. Table III illustrates the specific number of UE employed in the four discrete slices and the UE type was sourced from the open dataset [40]. Utilizing the above data, we have compiled a summary of the quantity of mobile UE presented during each hour and have subsequently allocated their respective start positions randomly within our 5G network. Fig 6 depicts the testing environment which is capable of accommodating diverse scenarios, including low and high peak traffic for both downlink and uplink directions (slice 3 - mMTC), stable traffic (slice 4 - VoIP), downlink direction (slice 1 - eMBB) exclusively, and uplink direction (slice 2 - uRLLC) exclusively.
V-B Proactive closed-loop algorithm - PCLANSA
The closed-loop algorithm is a crucial component of the service assurance mechanisms in 5G and B5G networks, and it is designed to oversee and enhance the performance of multiple elements of the network, and to ensure there are enough resources to satisfy the network SLAs. This section will evaluate PCLANSA’s performance in the service assurance domain within a 5G network. PCLANSA was designed with the intention of providing flexible configurations, to align with the requirements of infrastructure and network planning and to enable dynamic monitoring of network conditions, identify potential resource issues, and take appropriate actions to maintain high service quality. Thus, the outcome of the algorithm may vary based on the configuration parameters, and there exists a certain degree of compromise between over-provisioning and the number of network actions executed. For the sake of simplicity and to easily observe the results of our experiments, we use the same settings for all slices. However, in practice, it is possible to configure different parameters for each individual slice.
@@ V-B Proactive closed-loop algorithm - PCLANSA
PCLANSA was extensively examined to assess its effectiveness in two distinct layers: CN and TN layers. The evaluation process included various factors, including latency, throughput, jitter, and packet loss to maximize user experience. In order to conduct a comprehensive analysis of the algorithm’s capabilities, a set of E2E KPI limits and scale factors on different slices was established. Detailed information about these KPI limits can be found in Table IV. These factors play an important role in provisioning optimal service in 5G/B5G networks and provide insights into the algorithm’s ability to detect anomalies, adapt to new network parameters, and make real-time adjustments to optimize service performance. By inspecting the algorithm’s performance with respect to latency (delay experienced during data transmission), we aim to determine if PCLANSA is capable of minimizing delays and enhancing the overall responsiveness of the network. Throughput, the rate of data transmission through the network, is also a crucial factor under consideration. It is an essential measure of the algorithm’s capacity to efficiently manage large data volumes and sustain a consistent information flow. Jitter, the irregular variation of packet arrival times, is also taken into account during the evaluation process. The ability of the algorithm to mitigate jitter is of extreme importance, as it directly impacts the stability and consistency of data transmission. By minimizing jitter, the algorithm can significantly enhance the reliability and smoothness of network performance. Packet loss, the loss of data packets during transmission, is another critical parameter that must be evaluated. The capacity of the algorithm to detect and minimize packet loss is vital for maintaining the integrity and completeness of the transmitted information. Through the reduction of packet loss, the algorithm ensures that data is reliably and accurately delivered, preventing any information loss or service disruption.
@@
KPI Type/ Unit KPI Slice type
@@
Delay ms Average Delay 300 0.20 10 2.5 30(i) 2 100 0.25
@@
Throughput Kbps Throughput N/A N/A N/A N/A N/A N/A N/A N/A
Th: Threshold; CF: Scale Factor; Source: ($i$) Table 3.1 [41]; ($ii$) Table 3 [42]; ($iii$) Tables 14 [43]; ($iv$) Table 3&10 [44]; ($v$) Section I [45]
@@
$\rho^{\textsc{op}}$15 .1 .05
$\rho^{\textsc{s}}$05 - -
$\varepsilon$814 - -
$\rho^{\textsc{ru}}$8 - -
$\rho^{\textsc{d}}$02 - -
@@
Setting Slice Number of KPI violation Total action Scale_down Scale_up Scale_out Scale_in Ratio action / simulation time Ratio action hourly
1 eMBB uRLLC mMTC VoIP 0 0 0 0 238 130 223 136 15 29 8 20 216 101 211 116 4 0 2 0 3 0 2 0 33% 18.2% 30% 19.7% 1.38% .76% 1.25% .86%
2 eMBB uRLLC mMTC VoIP 9 (delays) 0 0 0 243 154 65 245 13 33 20 63 219 121 232 182 6 0 7 0 5 0 6 0 33.9% 21.5% 37% 35.3% 1.41% .9% 1.54% 1.47%
3 eMBB uRLLC mMTC VoIP 24 (delays), 2 (packet losses) 0 1 (delay) 0 349 168 321 294 104 45 50 101 234 123 258 193 6 0 7 0 5 0 6 0 48.7% 23.5% 44.8% 42.3% 2.03% .99% 1.86% 1.76%
@@
Figure 8: The utilization of VNFs allocated to the eMBB slice with the setting 1. (vertical green line: time of adding VNF, vertical red line: time of removing VNF, orange line: actual resource utilization, blue line: configure resource.
Table V outlines the configuration parameters utilized in the testing environment. During the testing phase, the algorithm successfully determined the data rate needed to be used as a traffic boundary for each slice, even when the slice exhibited a significantly high data rate. By taking the advanced ML traffic forecasting model, PCLANSA reliably offers the guaranteed bit rate and ensures a seamless flow of traffic.As an example refer to Fig. 7(a) for more details of the output configuration on the eMBB slice. On the other hand, the algorithm exhibits proficiency in efficiently providing an optimized allocation of resources of the VNF instances which can see in Fig. 8. The orange colour represents the actual resource utilization, while the blue colour represents the configured resources for the VNF instances. It can be observed that the algorithm has the capability to predict resource utilization and proactively allocate resources. As illustrated in Fig. 8(a) for the overall of all VNF instances of the eMBB slice, the algorithm closely provides optimal resources for the slice and always allocates spare resources in advance in accordance with the requirements of the slice. To elaborate on Fig. 8(b to f), the algorithm is capable of providing the necessary resources for each VNF instance while being able to dynamically add or remove instances correctly to optimize resources in accordance with the requirements of the slice. Therefore, this indicates that the algorithm offers the capability to identify and allocate resources in an optimized way effectively. Nevertheless, it is important to note that the distribution of spare resources and the execution of actions of the algorithm might vary based on the parameters $\rho^{\textsc{op}},\rho^{\textsc{ru}},\rho^{\textsc{d}}$ and $\rho^{\textsc{s}}$. If their values are sufficiently small, the algorithm will probably execute actions more frequently, as the available resources will be depleted sooner but the network will save more resources. In addition, the algorithm’s ability to effectively utilize resources can enhance the dependability and scalability of VNF deployments. Thus, the algorithm has the potential to mitigate bottlenecks and outages by promoting the timely and efficient utilization of resources. Furthermore, this can enhance the resilience of VNF deployments in the face of disruptions and improve their ability to handle high volumes of network traffic.
@@
The results of the algorithm on the simulation are shown in Table VI and Fig. 9. The closed-loop algorithm is effective at minimizing the number of KPI violations, even in the face of high traffic spikes and changes in network conditions while balancing the trade-off between the number of actions taken (around 0.86% to 1.38% hourly on setting 1) and the spare resources allocated to the slice. Depicts in Fig. 9, the algorithm with setting 1 successfully prevents violations of KPIs for all slices in terms of packet loss (a,b), delay(c,d), and jitter (e,f), thereby meeting our QoS targets. The simulations show that the closed-loop algorithm is capable of maintaining the QoS of multiple slices. It rapidly and accurately identified KPI violations, implemented corrective measures, took corrective actions, and allocated suitable resources to resolve KPI violations in a timely manner. Consequently, the algorithm was able to reduce the number of KPI violations over time, enhancing the overall QoS of the network.
@@
Assuming that the highest peak of network traffic is known and enough resources are configured for the slice to run smoothly during the simulation without any KPI violation, as an example in Fig. 10, the green colour represents the spare resources, and the orange colour represents the actual resources consumption of eMBB in the above case without our closed-loop algorithm. Our closed-loop algorithm can save 54.85% of overall resources in comparison to the case run without PCLANSA (see Fig. 8(a) and Fig. 10). The overall resource savings is the mean difference between the total resources used without the algorithm and the total resources used with the algorithm enabled in a given slice over the simulation time. Furthermore, the algorithm can achieve significant overall resource savings of 50.87% for mMTC, 57.1% for uRLLC, and 23.63% for VoIP. The relatively lower overall resource savings for the VoIP slice can be understood due to its stable traffic, as described in Section V-A. An inverse correlation between the accepted over-provisioning rate and the quantity of scaling actions was observed across various settings. When the accepted over-provisioning rate increases, the number of scaling actions decreases, whereas a reduction in the accepted over-provisioning rate results in an increase in scaling actions. This correlation can be understood by considering the trade-off between resource utilization and service assurance. When the accepted over-provisioning rate is high, the algorithm allows for a larger buffer of resources to be allocated to VNFs. This buffer can help to mitigate oscillations in demand without the need for additional scaling actions, resulting in a reduction in the number of scaling actions. Conversely, when the accepted over-provisioning rate is low, the algorithm allocates fewer resources to VNFs, leaving less room for absorbing rapid fluctuations in demand. As a result, the algorithm may need to take more frequent scaling actions in order to maintain service assurance, leading to an increase in the number of scaling actions. On the other hand, the closed-loop algorithm is also able to successfully scale resources in response to changes in traffic load, without experiencing any performance bottlenecks or outages in both horizontal, Fig. 7(b), and vertical scaling, Fig 8. In particular, it was able to accurately predict future resource demand and allocate resources. This allowed the algorithm to minimize resource usage while still meeting service assurance requirements. Additionally, because the algorithm is able to perform parallel tasks on different slices, it is able to simultaneously allocate resources to VNF instances on multiple slices. This allowed the algorithm to effectively manage resource allocation across all slices in real-time. For example, when there was a sudden increase in traffic on one slice, the algorithm was able to quickly allocate additional resources to that slice in order to avoid any KPI violation. At the same time, if the traffic decreased on another slice, the algorithm was able to reduce resource allocation accordingly to minimize resource usage.
@@ VI Conclusion
The proactive closed-loop algorithm for service assurance described in this paper, PCLANSA, represents a promising approach to effectively managing resource allocation in 5G/B5G networks. By combining ML and linear programming techniques, the algorithm is able to accurately predict future resource demand and allocate resources accordingly. An important advantage of the algorithm consists of its ability to effectively manage the trade-off between the number of actions performed and the allocation of spare resources to the network slice. This allows the algorithm to minimize resource usage while still meeting service assurance requirements in different slices. Another important strength of the algorithm is its ability to simultaneously allocate resources to VNFs on multiple slices, taking into account the specific service assurance requirements and resource demands of each slice. Additionally, it has the ability to perform in parallel on different slices allowing it to effectively manage resource allocation across all slices in real-time with scalability. Our results demonstrate the reliability, robustness and efficacy of PCLANSA in managing resource allocation across multiple network slices. In detail, PCLANSA has demonstrated significant resource savings across all slices, with up to 54.85% savings in the eMBB slice, 50.67% in the mMTC slice, 57.1% in the uRLLC slice, and 23.63% in the VoIP slice. Even when decreasing the spare resources configuration to a small ratio (5%), the algorithm remains highly effective, with very few KPI violations. Overall, our proposed proactive closed-loop algorithm for service assurance represents a significant advance in the management of resource allocation in the network and has the potential to significantly improve the efficiency and effectiveness of 5G/B5G networks.

### v4

@@ V Evaluation
This section will provide an overview of our experimental setup and showcase our service assurance algorithm designed to support 5G networks with network slices, encompassing diverse network slice categories.
V-A 5G network slice environment
A packet-level simulation was developed using Omnet++ [30] to emulate a 5G network environment, incorporating support for slicing features. The 5G E2E network slice simulation involves the initial configuration of four distinct network slices: uRLLC (video gaming), mMTC (IoT), eMBB (HD video), and an intermediary application service, such as VoIP. Each slice is designed to meet unique service assurance requirements and resource demands. Fig. 5 demonstrates the implementation of our 5G network, which is reinforced by an isolated E2E network slice mechanism that leverages virtualization technology. The 5G CN enables the construction of VNFs in a dynamic manner, as displayed by the User Plane Function (UPF) in our experimentation. This enables a single network slice to accommodate either a singular VNF or a group of VNFs supported by a load balancer. Hence, the CN has the capability of facilitating the scaling of VNFs in both the vertical and horizontal dimensions. To balance traffic between VNF instances within a network slice, a weighted round-robin load balancing algorithm [31] was integrated into the network slice manager. In order to address the guaranteed bit rate requirements in network slicing, a Hierarchical Token Bucket (HTB) queue [32] has been implemented in the router. This queue has been shown to assist in isolating virtual network links in both TN and CN, thereby optimizing resources.
@@
network slice Transport type Total Scale factor Network direction
@@
TABLE III: Mapping from open data to the network slice devices.
@@
Our simulation was configured to generate traffic patterns that closely resemble those found in real-world networks. Table III illustrates the specific number of UEs used in the four distinct network slices, with the UE types sourced from the open dataset [33]. Using the data mentioned above, we compiled a summary of the number of mobile UEs present during each hour and randomly allocated their respective start positions within the 5G network. Fig. 6 depicts the testing environment, which is capable of accommodating diverse scenarios, including low and high peak traffic for both downlink and uplink directions (network slice 3 - mMTC), stable traffic (network slice 4 - VoIP), downlink direction (network slice 1 - eMBB) exclusively, and uplink direction (network slice 2 - uRLLC) exclusively.
V-B Evaluation of PCLANSA
This section will evaluate PCLANSA’s performance in the service assurance domain within a 5G network environment. The evaluation will be conducted within the simulation environment described above. Our PCLANSA was designed to offer flexible configurations, aligning with infrastructure and network planning requirements. It enables dynamic monitoring of network conditions, helping to identify potential resource issues and take appropriate actions to maintain high service quality for the network slice. In the interest of simplicity and the capacity to readily discern the outcomes of our experiments, we employ uniform settings for all network slices. Nevertheless, it is feasible to configure disparate parameters for each network slice in practice.
@@ V-B Evaluation of PCLANSA
PCLANSA was extensively examined to assess its effectiveness in two distinct layers: CN and TN layers. The evaluation process included various factors, including latency, throughput, jitter, and packet loss, to maximize user experience. To conduct a comprehensive analysis of the algorithm’s capabilities, a set of E2E KPI limits and scale factors on different network slices was established and collected from different references [34, 35, 36, 37, 38]. Detailed information about these KPI limits can be found in Table IV. These factors play an important role in provisioning optimal service in 5G/B5G networks and provide insights into the algorithm’s ability to detect anomalies, adapt to new network parameters, and make real-time adjustments to optimize service performance.
@@
KPI Type/
Unit
KPI Network slice type
@@
Delay
ms
Average Delay 300 0.20 10 2.5 30(i) 2 100 0.25
@@
Throughput
Kbps
Throughput N/A N/A N/A N/A N/A N/A N/A N/A
Th: Threshold; CF: Scale Factor; Source: (ii) Table 3.1 [34]; (iiii) Table 3 [35]; (iiiiii) Tables 14 [36]; (iviv) Table 3&10 [37]; ($v$) Section I [38]
@@
$\rho^{\textsc{op}}$ .15 .1 .05
$\rho^{\textsc{s}}$ .05 - -
$\varepsilon$ .814 - -
$\rho^{\textsc{ru}}$ .8 - -
$\rho^{\textsc{d}}$ .02 - -
@@
Setting
Network
slice
Number of
KPI violation
Total
action
Scale_down Scale_up Scale_out Scale_in
Ratio action
/ simulation time
Ratio action
hourly
1
eMBB
uRLLC
mMTC
VoIP
0
0
0
0
238
130
223
136
15
29
8
20
216
101
211
116
4
0
2
0
3
0
2
0
33%
18.2%
30%
19.7%
1.38%
.76%
1.25%
.86%
2
eMBB
uRLLC
mMTC
VoIP
9 (delays)
0
0
0
243
154
65
245
13
33
20
63
219
121
232
182
6
0
7
0
5
0
6
0
33.9%
21.5%
37%
35.3%
1.41%
.9%
1.54%
1.47%
3
eMBB
uRLLC
mMTC
VoIP
24 (delays), 2 (packet losses)
0
1 (delay)
0
349
168
321
294
104
45
50
101
234
123
258
193
6
0
7
0
5
0
6
0
48.7%
23.5%
44.8%
42.3%
2.03%
.99%
1.86%
1.76%
@@
Figure 8: The utilization of VNFs allocated to the eMBB slice with setting 1. (vertical green line: time of adding VNF, vertical red line: time of removing VNF, orange line: actual resource utilization, blue line: configured resource.)
The configuration parameters utilized in the evaluation environment are delineated in Table V. During the evaluation phase, the algorithm successfully determined the data rate needed to be used as a traffic boundary for each network slice, even when the network slice exhibited a significantly high data rate. By leveraging an advanced ML traffic forecasting model, PCLANSA reliably guarantees the bit rate and ensures a seamless flow of traffic. For a more detailed view of the output configuration, refer to Fig. 7 (a) as an example of the eMBB slice. Meanwhile, the algorithm demonstrates its proficiency in optimizing the allocation of resources for the VNF instances, as depicted in Fig. 8. In the figure, the orange colour represents the actual resource utilization, while the blue colour indicates the configured resources for the VNF instances. It is evident that the algorithm excels in predicting resource utilization and proactively allocating resources accordingly. As illustrated in Fig. 8(a) for the overall utilization of all VNF instances of the eMBB slice, the algorithm closely provides optimal resources for the network slice and always allocates spare resources in advance in accordance with the requirements of the slice. To elaborate on Fig. 8(b to f), the algorithm is capable of providing the necessary resources for each VNF instance while being able to dynamically add or remove instances correctly to optimize resources in accordance with the requirements of the slice. Therefore, this indicates that the algorithm offers the capability to identify and allocate resources in an optimized way. Nevertheless, it is important to note that the distribution of spare resources and the execution of actions of the algorithm might vary based on the parameters $\rho^{\textsc{op}},\rho^{\textsc{ru}},\rho^{\textsc{d}}$ and $\rho^{\textsc{s}}$. If their values are sufficiently small, the algorithm will probably execute actions more frequently, as the available resources will be depleted sooner, but the network will save more resources.
@@
A comprehensive overview of the PCLANSA performance across four network slices is provided in both Table VI and Fig. 9. In detail, PCLANSA proves effective at minimizing KPI violations across various settings, as presented in Table V. This is achieved even during high traffic spikes and network condition changes, as demonstrated in the use cases of the eMBB and uRLLC slices. Furthermore, it strikes a balance between the number of actions taken (ranging from 1% to 2% on an hourly basis) and the allocation of spare resources to the network slices. As depicted in Fig. 9, the algorithm with setting 1 successfully prevents KPI violations across all network slices, in terms of packet loss (a, b), delay (c, d), and jitter (e, f), thus meeting our QoS targets. In addition, the results also demonstrate the efficacy of PCLANSA in executing parallel operations with a high level of performance. It quickly and accurately identifies KPI violations, performs corrective actions, and allocates appropriate resources to resolve issues promptly. Therefore, the PCLANSA effectively reduces the number of KPI violations over time, leading to improved overall QoS for the network and enhanced QoE for end users.
@@
Assuming that the highest peak of network traffic is known and sufficient resources are configured for a network slice to operate efficiently without any KPI violation. As illustrated in Figure 10, spare resources are represented in green, while actual eMBB resource consumption is depicted in orange. In comparison to this worst-case scenario, our PCLANSA can reduce resource consumption by 54.85% for the eMBB slice (see Fig. 8(a) and Fig. 10). The overall resource savings are calculated as the mean difference between the total resources used with and without the algorithm enabled, over the entire simulation period. Across disparate network slices, the algorithm has been shown to achieve significant aggregate resource savings. Specifically, resource savings of 50.87% for mMTC, 57.1% for uRLLC, and 23.63% for VoIP were observed. The relatively lower savings for VoIP can be attributed to its stable traffic, as discussed in Section V-A. Additionally, we observed an inverse correlation between the accepted over-provisioning rate and the number of scaling actions. As the over-provisioning rate increases, the number of scaling actions decreases. This is due to the trade-off between resource utilization and service assurance: higher over-provisioning provides a buffer that mitigates demand fluctuations, reducing the need for scaling actions. Conversely, a lower over-provisioning rate requires more frequent scaling actions to maintain service assurance, leading to an increase in scaling actions. Moreover, our closed-loop algorithm efficiently scales resources in response to traffic load changes without performance bottlenecks or outages, both in horizontal scaling (Fig. 7(b)) and vertical scaling (Fig. 8). It accurately predicts future resource demands, enabling optimal resource allocation while maintaining service assurance. The algorithm also handles concurrent tasks across different network slices by deploying multiple PCLANSA instances, coordinated through the Broker and Slice Control components. This real-time coordination allows the algorithm to dynamically allocate resources to VNF instances across multiple network slices. For example, in response to a traffic spike on one network slice, the algorithm quickly allocates additional resources, preventing KPI violations. Simultaneously, when traffic decreases on another network slice, it reduces resource allocation, optimizing resource usage.
@@ VI Conclusion
This paper presented PCLANSA, a proactive closed-loop algorithm for service assurance in 5G/B5G network slicing. PCLANSA dynamically scales VNF resources and manages link capacity to meet network slice-specific KPIs while minimizing resource consumption. By leveraging machine learning for traffic prediction and linear programming for resource optimization, PCLANSA proactively adapts to changing network conditions and prevents KPI violations. Our experimental results demonstrate significant resource savings across diverse network slice types. PCLANSA achieved up to 54.85%, 50.87%, 57.1%, and 23.63% resource savings for eMBB, mMTC, uRLLC, and VoIP slices, respectively, in comparison to the worst-case scenario. Even with minimal over-provisioning at just 5%, the PCLANSA algorithm remains highly effective, resulting in minimal KPI violations, with only 27 violations recorded over 24 hours of simulation. These results highlight PCLANSA’s potential to significantly improve the efficiency and effectiveness of resource management in 5G/B5G networks.
In future studies, we intend to further explore and develop the capabilities of PCLANSA to manage dynamic network slice creation and deletion. The objective is to integrate more sophisticated traffic prediction models with the incorporation of network topology. Furthermore, the evaluation of PCLANSA is planned to be conducted in a real-world testbed.